# Theoretical Investigation on Inductive Bias of Isolation Forest

### v1

@@ 4 Case studies for inductive bias
Definition 4.1 (Density Metrics).
Let $U\triangleq\max_{i\geq 2}~{}\lvert x_{i+1}-x_{i}\rvert$ and $L\triangleq\min_{i\geq 2}~{}\lvert x_{i+1}-x_{i}\rvert$ be the maximum and minimum distances between adjacent points, respectively. Define two density metrics:
  • The density factor $\kappa$ measures the ratio of $U$ to $L$: $\kappa=U/L$.
  • The density difference $\delta$ measures the difference between $U$ and $L$: $\delta=U-L$.
These metrics evaluate how uniformly data points are distributed relative to their neighbors. Intuitively, a dataset composed entirely of normal points is expected to exhibit small values for both metrics; otherwise, some points would be excessively distant from their neighbors. In the case studies below, we demonstrate that these density metrics play a critical role in shaping the inductive bias of iForest.
Before proceeding, we introduce the following assumption.
@@
Assume that density factor $\kappa\geq\sqrt{n+3}$.
Assumption 4.2 is frequently observed in practice. To justify its reasonability, we check all dimensions of binary classification datasets from the OpenML benchmark [Vanschoren et al., 2014, Bischl et al., 2021]. The results are summarized in Table 1. “Successful” indicates that the density factor exceeds $\sqrt{n+3}$. “Valid” indicates the nonexistence of repeated values in the dimension. Notably, the theoretical probability of repeated values is zero. Thus, we omit this case for the convenience of analysis. Our analysis demonstrates that the density factor exceeds $\sqrt{n+3}$ in almost all dimensions across all datasets, highlighting the mildness of Assumption 4.2. A theoretical study of Assumption 4.2 is also conducted, with detailed results available in Appendix B.
Table 1: Number of features in OpenML benchmark that verify Assumption 4.2.
@@ 4.1 Marginal single anomaly
The first case study is about the marginal single anomaly, an example of which is shown in Figure 2 (a), where the anomaly is located at the leftmost and the normal points are distributed at the right side. Marginal single anomalies may be the most common ones in practice, which happens when there is some noise in normal data drawn from continuous and bounded support. Note that although this model looks simple, it is the basic building block of more complex models.
@@
Theorem 4.3 shows that the expected depth of the marginal single anomaly is the smallest among all the data points when $x_{2}-x_{1}>U\cdot\kappa$. Hence, the marginal single anomaly is the most likely to be the anomaly output by iForest. For the rightmost marginal single anomaly Theorem 4.3 holds similarly. Theorem 4.3 shows the sufficiency of $x_{2}-x_{1}>U\cdot\kappa$. The following theorem shows the necessity.
@@ 4.2 Central single anomaly
Central single anomaly is another common type of anomaly, as shown in Figure 2 (b). The anomaly is located at the center of the data points and the normal points are distributed at both the left and right sides. This models the scenario when the data distribution consists of multiple clusters, and the anomaly is likely to arise in the gap between different clusters.
@@ 4.3 Marginal clustered anomalies
In this subsection, we consider the case of marginal clustered anomalies. The motivation is that anomaly events may occur multiple times, as shown in Figure 2 (c). A group of anomalies is located at the leftmost of the data points, and the normal points are distributed on the right side. Marginal clustered anomalies characterize some unimodal data with multiple clustered noisy points.
@@
Let $n_{1}$ be an odd number for convenience. Suppose $n_{1}=o(n_{0})$. For any dataset $D$, of which $x_{1:n_{1}}$ and $x_{(n_{1}+1):(n_{1}+n_{0})}$ verify $\kappa$-dense, the sufficient and necessary condition for iForest to detect all the marginal clustered anomalies is that $x_{n_{1}+1}-x_{n_{1}}>\Theta(n_{1}^{2}\kappa)$.
@@
Table 2 summarizes three cases of anomalies in this section. It is observed that iForest acts to be more cautious in reporting an anomaly as the threshold is relatively large but more parameter-adaptive to different types of anomalies than $k$-NN as the thresholds of iForest are only problem-dependent.
Table 2: Decision thresholds of iForest (IF) and $k$-NN for different types of anomalies.
Anomaly type IF $k$-NN
@@
(b) HTTP dataset
(c) SMTP dataset
Figure 3: Mean-square errors of the real depths learned by iForest to the theoretical expected depths about numbers of trees. The shaded regions represent 95% confidence regions over multiple runs.
@@ 5.1 Convergence of the empirical depth
This experiment is conducted using both synthetic and real-world datasets. We begin by independently sampling 100 points from three distributions: the standard normal distribution $\mathcal{N}(0,1)$, the uniform distribution $U[0,1]$, and the exponential distribution $\textrm{Exp}(1)$. Additionally, we examine the HTTP and SMTP datasets [UCI KDD Archive, 1999], which are widely used anomaly detection datasets. Focusing on the one-dimensional case, we separately select each feature from all datasets to demonstrate the convergence of the empirical depth. For each dataset, we evaluate the mean squared errors of the depths learned by iForest to the theoretical expected depths with 100 subsamples as the number of iTrees ranges from 100 to 1000. Each setup is repeated 10 times independently.
Figure 3 shows the curves of the errors about the number of iTrees. The error decreases rapidly as the number of trees increases, which implies that the depth of every point may eventually converge to the theoretically expected depth. This verifies the concentration property of the empirical depth and the correctness of the expected depth function shown in Prososition 3.1 and Theorem 3.5, respectively.
@@ 5.2 Anomaly detection of iForest
We generate multiple datasets, each of which contains 20 points drawn from a uniform distribution independently. We compute the expected depth of each point to help understand the behavior of iForest. The results are presented in Figure 4, where we label the points with a small depth considered as anomalies and a larger depth considered as normal points in red and blue colors, respectively.
@@ 6 Discussions
The above analysis focuses on one-dimensional cases. In this section, we extend to multi-dimensional cases, including reducible and irreducible cases.
(a) Reducible case 1
(b) Reducible case 2
(c) Irreducible case
Figure 5: The expected depth function in the reducible and irreducible cases.
Reducible case. We say that the anomalies are reducible if they remain anomalous when projected on any dimension. For example, Figures 5 (a) and (b) can be reduced to marginal and central anomalies when projected on both dimensions, respectively. In this case, we can estimate the expected depth by the average depth projected on each dimension, defined as
$\bar{h}(\mathbf{x};\mathbf{x}_{1},\dots,\mathbf{x}_{n})\approx\frac{1}{d}\sum_{j=1}^{d}\bar{h}(\mathbf{x}^{(j)};\mathbf{x}_{1}^{(j)},\dots,\mathbf{x}_{n}^{(j)})\ .$
Irreducible case. If the anomalies are irreducible, we can first apply some feature mapping $\phi:\mathbb{R}^{d}\to\mathbb{R}$ and estimate the expected depth by the depth of $\phi(\mathbf{x})$, defined as
$\bar{h}(\mathbf{x};\mathbf{x}_{1},\dots,\mathbf{x}_{n})\approx\bar{h}(\phi(\mathbf{x});\phi(\mathbf{x}_{1}),\dots,\phi(\mathbf{x}_{n}))\ .$
If the feature mapping $\phi$ can suitably model the data distribution, one can obtain an ideal estimate of the expected depth. For example, a radial basis function (RBF) kernel mapping perfectly makes the data highly separable, as shown in Figure 5 (c). This also reduces to marginal single anomalies.
Through direct reduction or kernel mapping, we can reduce the multi-dimensional anomaly detection to one-dimensional cases and further extend our one-dimensional analysis. Hence, all the case studies and conclusions in Section 4 can be applied to the multi-dimensional cases.
@@ Appendix A Proofs in Section 3
Proof   We first denote $h^{\prime}_{m}=M^{-1}h_{m}(\mathbf{x};\mathbf{\Theta}_{m})$. From the growth process of Isolation Trees, we have that $h^{\prime}_{m}$ verifies i.i.d. condition, $h^{\prime}_{m}\in(0,n/M)$, and
$\mathbb{E}_{\mathbf{\Theta}}[h^{\prime}_{m}]=\mathbb{E}_{\mathbf{\Theta}}[h(\mathbf{x};\mathbf{\Theta})],~{}\forall 1\leq m\leq M\ .$
By applying Hoeffding’s inequality (Lemma A.1) to $h^{\prime}_{m}$, we have
$\Pr\left[\left|\sum_{m=1}^{M}h^{\prime}_{m}-\mathbb{E}\left[\sum_{m=1}^{M}h^{\prime}_{m}\right]\right|\geq\epsilon\right]$ $\leq 2\exp\left(-2\epsilon^{2}/(M(n/M)^{2})\right)$
@@
Transition probability. We now verify the transition probability. Recall that the current state, the next state, and the absorbed state are denoted by $(x_{\ell},x_{r})$ and $(x_{\ell^{\prime}},x_{r^{\prime}})$, $(x_{i},x_{i})$, respectively, where $x_{\ell},x_{\ell^{\prime}},x_{r},x_{r^{\prime}},x_{i}\in D$, $x_{\ell}\leq x_{i}\leq x_{r}$, and $x_{\ell^{\prime}}\leq x_{i}\leq x_{r^{\prime}}$. In every step, Isolation Tree chooses the split value in the interval $(x_{\ell},x_{i})$ or interval $(x_{i},x_{r})$, and thus either $x_{\ell^{\prime}}=x_{\ell}$ or $x_{r^{\prime}}=x_{r}$ holds, implying that
@@
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]\ .$
Let $\mathbf{\Theta}=(\Theta_{1},\dots,\Theta_{\tau})$ be the vector of random variables representing the split at each step, where $\tau$ denotes the depth of point $x_{i}$. Note that $\tau$ can be indeterminate, as the depth is a random variable. Let $\Omega$ be the event that the first split is between $x^{\prime}$ and $x_{1}$. Then, we can expand the expectation as follows:
$\mathbb{E}_{\mathbf{\Theta}}\Big{[}h\big{(}x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\big{)}\Big{]}$ $=\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[\mathbb{E}_{\Theta_{1}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]\right]$
$=\Pr_{\Theta_{1}}\left[\Theta_{1}\in\Omega\right]~{}\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\in\Omega\right]$
$\quad\Pr_{\Theta_{1}}\left[\Theta_{1}\notin\Omega\right]~{}\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\notin\Omega\right]$
$=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}~{}\left(1+\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\in\Omega\right]\right)+$
$\quad\frac{x_{i}-x_{1}}{x_{i}-x^{\prime}}~{}\left(1+\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\notin\Omega\right]\right)\ .$
For $\Theta_{1}\in\Omega$, the tree growth process after the first split is the same as the case where $x^{\prime}$ is not added
$\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\in\Omega\right]=\mathbb{E}_{\mathbf{\Theta}^{\prime}_{1:\tau}}\left[h\left(x_{i};x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]\ .$
For $\Theta_{1}\notin\Omega$, in fact, the tree growth process is the same as the case where $x^{\prime}$ is added, i.e.,
$1+\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\notin\Omega\right]=\mathbb{E}_{\mathbf{\Theta}^{\prime}_{1:\tau}}\left[h\left(x_{i};x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]\ .$
@@
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]$ $=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}~{}\left(1+\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\in\Omega\right]\right)+$
$\quad\frac{x_{i}-x_{1}}{x_{i}-x^{\prime}}~{}\left(1+\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\notin\Omega\right]\right)$
@@
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right]=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right]\ .$ (2)

### v2

@@ 4 INDUCTIVE BIAS
Definition 4.1.
Let $U\triangleq\max_{i\geq 2}~\lvert x_{i+1}-x_{i}\rvert$ and $L\triangleq\min_{i\geq 2}~\lvert x_{i+1}-x_{i}\rvert$ be the maximum and minimum distances between adjacent points, respectively. Define two density metrics as follows:
  • Density ratio: $\kappa=U/L$.
  • Density difference: $\delta=U-L$.
The two metrics evaluate how non-uniformly data are distributed relative to their neighbors, measured by ratio and difference. Intuitively, a dataset composed entirely of normal points is expected to exhibit small values for both metrics; otherwise, some points would be excessively distant from their neighbors. In the case studies below, we show that these density metrics play a critical role in shaping the inductive bias of iForest.
Before proceeding, we introduce the assumption.
@@
Assume that $\kappa\geq\sqrt{n+3}$.
Assumption 4.2 is frequently observed in practice. To justify its reasonability, we check all dimensions of binary classification datasets from the OpenML benchmark [Vanschoren et al., 2014, Bischl et al., 2021]. The results are summarized in Table 1. “Successful” indicates that the density ratio exceeds $\sqrt{n+3}$. “Valid” indicates the nonexistence of repeated values in the dimension. Notably, the theoretical probability of repeated values is zero. Thus, we omit this case for the convenience of analysis. Our analysis demonstrates that the density ratio exceeds $\sqrt{n+3}$ in almost all dimensions across all datasets, highlighting the mildness of Assumption 4.2. A theoretical study of the mildness is also conducted, with detailed results in Appendix B.
Table 1: Number of features in OpenML benchmark that successfully verify Assumption 4.2.
@@ 4.1 Marginal single anomaly
The first case study is about the marginal single anomaly, an example of which is shown in Figure 2, where the anomaly is located at the leftmost, and the normal points are distributed on the right side. For the right-most anomaly with normal points on the left, we can have a similar conclusion. Marginal single anomalies may be the most common ones in practice, which happens when there is some abrupt change in normal data drawn from continuous and bounded support. Note that although this model looks simple, it is the basic building block of more complex models.
Figure 2: Marginal single anomaly.
@@
Theorem 4.3 shows that the expected depth of the marginal single anomaly is the smallest among all the data points when $x_{2}-x_{1}>U\cdot\kappa$. Hence, the marginal single anomaly is the most likely to be the anomaly output by iForest. For the rightmost marginal single anomaly, Theorem 4.3 holds similarly. Theorem 4.3 shows the sufficiency of $x_{2}-x_{1}>U\cdot\kappa$. The following theorem shows the necessity.
@@ 4.2 Central single anomaly
Central single anomaly is another common type of anomaly, as shown in Figure 3. The anomaly is located at the center of the data points, and the normal points are distributed on both the left and right sides. This models the scenario when the data distribution consists of multiple clusters, and the anomaly is likely to arise in the gap between different clusters.
Figure 3: Central single anomaly.
@@ 4.3 Marginal clustered anomalies
In this subsection, we consider the case of marginal clustered anomalies. The motivation is that anomaly events may occur multiple times, as shown in Figure 4. A group of anomalies is located at the leftmost, and the normal points are distributed on the right. Marginal clustered anomalies characterize some unimodal data with multiple clustered noisy points.
Figure 4: Marginal multiple anomalies.
@@
Let $n_{1}$ be an odd number for convenience where $n_{1}=o(n_{0})$. For any dataset $D$, of which $x_{1:n_{1}}$ and $x_{(n_{1}+1):(n_{1}+n_{0})}$ verify $\kappa$-dense, the sufficient and necessary condition for iForest to detect all the marginal clustered anomalies is $x_{n_{1}+1}-x_{n_{1}}>\Theta(n_{1}^{2}\kappa)$.
@@
Table 2: Decision thresholds of iForest and $k$-NN for different types of anomalies.
Anomaly type iForest $k$-NN
@@
Table 2 summarizes three cases of anomalies in this section. It is observed that iForest acts to be more cautious in reporting an anomaly as the threshold is relatively large, but more parameter-adaptive to different types of anomalies than $k$-NN, as the thresholds of iForest are only problem-dependent.
@@
(b) Iris dataset
(c) HTTP dataset
(d) SMTP dataset
Figure 5: Mean-square errors of the real depths learned by iForest to the theoretical expected depths about numbers of trees. The shaded regions represent the confidence regions over multiple runs.
@@ 5.1 Convergence of the empirical depth
This experiment is conducted using both synthetic and real-world datasets. We begin by independently sampling 100 points from three distributions: the standard normal distribution $\mathcal{N}(0,1)$, the uniform distribution $U[0,1]$, and the exponential distribution $\textrm{Exp}(1)$. For real-world datasets, we examine the iris dataset [Fisher, 1936], which is a widely used machine learning dataset. Additionally, we examine the HTTP and SMTP datasets [UCI KDD Archive, 1999], which are widely used anomaly detection datasets. Focusing on the one-dimensional case, we separately select each feature from all datasets to demonstrate the convergence of the empirical depth. For each dataset, we evaluate the mean squared errors of the depths learned by iForest to the theoretical expected depths with 100 subsamples as the number of iTrees ranges from 100 to 1000. Each setup is repeated 10 times independently.
Figure 5 shows the curves of the errors about the number of iTrees for different datasets. We can see that the error decreases rapidly as the number of trees increases, which implies that the depth of every point may eventually converge to the theoretically expected depth. This verifies the concentration property of the empirical depth to the theoretical depth and the correctness of our derived expected depth function shown in Proposition 3.1 and Theorem 3.5, respectively.
@@ 5.2 Anomaly detection of iForest
We generate multiple datasets, each of which contains 20 points drawn from a uniform distribution independently. We compute the expected depth of each point to help understand when and when not iForest can detect anomalies. The results are presented in Figure 6, where we label the points with a small depth considered as anomalies and a larger depth considered as normal points in red and blue colors, respectively.
@@ 6 DISCUSSIONS
Section 3 provided theoretical studies of iForest in the one-dimensional case. In this section, we proceed to discuss the extended results of multi-dimensional cases.
Challenges in Multi-Dimensional Extensions
In essence, there are two main challenges for theoretically analyzing the inductive bias of Isolation Trees in multi-dimensional cases. Firstly, there emerge a lot of relative positions between the concerned point and others for the multi-dimensional cases, which usually lead to various forms of the expected depth function. For example, in the simple two-dimensional cases with five points, there are at least six different forms of expected depth function for different point permutations, as shown in Figure 7. Thus, it is difficult to employ a unified scheme to handle various expected depth functions. Secondly, the candidate splitting attribute and splitting point may completely change after one split due to the randomness of splittings. Taking Figure 8 as an example, after one split, there are two candidate split attributes in the leftmost case but only one in the middle and right cases, which makes the growth process of Isolation Trees unpredictable. By applying the same random walk model, just like the one-dimensional cases, the state space grows exponentially with the dimension $d$, making the analysis intractable. And there are almost no mature techniques for random processes to learn from [Bhattacharya and Waymire, 2021].
A Negative Result in Two Dimensions
To formally demonstrate the difficulty, we studied a two-dimensional toy example with four nearly collinear points. We have the following negative result:
Proposition 6.1.
For $D=\{\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\mathbf{x}_{4}\}$ where $\mathbf{x}_{1}=(1,1)$, $\mathbf{x}_{2}=(2,3)$, $\mathbf{x}_{3}=(3,4)$, and $\mathbf{x}_{4}=(4,5)$, we have the followings:
$\frac{745}{288}=\bar{h}(\mathbf{x}_{3})$ $\neq\bar{h}(\mathbf{x}_{3};\mathbf{x}_{1:3})+\bar{h}(\mathbf{x}_{3};\mathbf{x}_{3:4})=\frac{744}{288}\ ,$
$\frac{491}{288}=\bar{h}(\mathbf{x}_{1})$ $\neq\frac{\bar{h}(\mathbf{x}_{1}^{(1)};\mathbf{x}_{1:4}^{(1)})+\bar{h}(\mathbf{x}_{1}^{(2)};\mathbf{x}_{1:4}^{(2)})}{2}=\frac{492}{288}\ ,$
where $\bar{h}(\mathbf{x}_{1};\mathbf{x}_{1:4})$ and $\bar{h}(\mathbf{x}_{3};\mathbf{x}_{1:4})$ are abbreviated as $\bar{h}(\mathbf{x}_{1})$ and $\bar{h}(\mathbf{x}_{3})$, respectively.
Proposition 6.1 shows that even in such a simple two-dimensional setting, basic characteristics fail to hold: (1) the expected depth cannot be decomposed into contributions from left and right subsets; (2) it cannot be reduced to the average over the projections on all dimensions. These violations indicate that closed-form expressions of the expected depth function may not exist in multi-dimensional settings and may require case-by-case analysis for different relative positions.
Although deriving exact formulas for multi-dimensional cases may be infeasible, it may be feasible to establish upper and lower bounds of normal and anomalous points in the purpose line of inductive bias analysis. Such analysis can be converted into identifying best- and worst-case scenarios, which are still tractable using the random walk framework, and it is interesting to explore this in the future.
In summary, while the theoretical analysis of iForest in multi-dimensional spaces remains challenging, our one-dimensional study provides essential foundations. The random walk model introduced here not only enables a precise characterization of the inductive bias in one dimension but also lays the groundwork for future explorations in multi-dimensional settings.
Figure 7: An example of six different points relative positions that lead to different forms of the expected depth function. Red points are the concerned points.
Figure 8: An example of a candidate split attribute and a split point changing after one split. Red points are the concerned points.
@@ 7 CONCLUSIONS
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model.
      Yes, Sections 2-3 provide the detailed description.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm.
      Not Applicable, this paper is a theoretical paper studying existing algorithms.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries.
      Yes, please refer to the supplemental material.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results.
      Yes, Sections 2-4 provide the detailed assumptions.
    2. (b)
      Complete proofs of all theoretical results.
      Yes, please refer to the Appendix.
    3. (c)
      Clear explanations of any assumptions.
      Yes, we have detailed explanations after all the assumptions.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL).
      Yes, please refer to the supplemental material.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen).
      Yes, please refer to the supplemental material.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times).
      Yes, please refer to the supplemental material.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider).
      Yes, please refer to the supplemental material.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets.
      Yes, we have cited the creator in the bibliography.
    2. (b)
      The license information of the assets, if applicable.
      Not Applicable.
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable.
      Not Applicable.
    4. (d)
      Information about consent from data providers/curators.
      Not Applicable.
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content.
      Not Applicable.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots.
      Not Applicable.
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable.
      Not Applicable.
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation.
      Not Applicable.
@@ Appendix A Proofs in Section 3
Proof We first denote $h^{\prime}(\mathbf{x};\mathbf{\Theta}_{m})=M^{-1}h(\mathbf{x};\mathbf{\Theta}_{m})$. From the growth process of Isolation Trees, $\mathbf{\Theta}_{1},\dots,\mathbf{\Theta}_{m}$ are i.i.d. copies of $\mathbf{\Theta}$, and we have
$\mathbb{E}_{\mathbf{\Theta}}[h^{\prime}(\mathbf{x};\mathbf{\Theta}_{m})]=\mathbb{E}_{\mathbf{\Theta}}[h^{\prime}(\mathbf{x};\mathbf{\Theta})],~\forall 1\leq m\leq M\ .$
By applying Hoeffding’s inequality (Lemma A.1) to $h^{\prime}(\mathbf{x};\mathbf{\Theta}_{m})$, we have
$\Pr\Bigg[\Bigg|\frac{1}{M}\sum_{m=1}^{M}h(\mathbf{x};D,\mathbf{\Theta}_{m})-\mathbb{E}_{\mathbf{\Theta}}[h(\mathbf{x};D,\mathbf{\Theta})]\Bigg|\geq\epsilon\Bigg]$ $=\Pr\left[\left|\sum_{m=1}^{M}h^{\prime}(\mathbf{x};\mathbf{\Theta}_{m})-\mathbb{E}\left[\sum_{m=1}^{M}h^{\prime}(\mathbf{x};\mathbf{\Theta}_{m})\right]\right|\geq\epsilon\right]$
$\leq 2\exp\left(-2\epsilon^{2}/(M(n/M)^{2})\right)$
@@
Transition probability. We now verify the transition probability. Recall that the current state, the next state, and the absorbed state are denoted by $(x_{\ell},x_{r})$ and $(x_{\ell^{\prime}},x_{r^{\prime}})$, $(x_{i},x_{i})$, respectively, where $_{\ell},x_{\ell^{\prime}},x_{r},x_{r^{\prime}},x_{i}\in D, xx$$_{\ell}\leq x_{i}\leq x_{r}$, and $x_{\ell^{\prime}}\leq x_{i}\leq x_{r^{\prime}}$. In every step, Isolation Tree chooses the split value in the interval $(x_{\ell},x_{i})$ or interval $(x_{i},x_{r})$, and thus either $x_{\ell^{\prime}}=x_{\ell}$ or $x_{r^{\prime}}=x_{r}$ holds, implying that
@@
$\mathbb{E}_{\mathbf{\Theta}^{\prime}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}^{\prime}\right)\right]=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]\ .$
Let $\mathbf{\Theta}^{\prime}=(\Theta^{\prime}_{1},\dots,\Theta^{\prime}_{i})$ and $\mathbf{\Theta}=(\Theta_{1},\dots,\Theta_{i-1})$ be the vectors of random variables representing the split at each step after $x^{\prime}$ is added and before $x^{\prime}$ is added, respectively. Note that when $x_{i}$ is isolated, the randomness can be determined. Let $\Omega$ be the event that the first split is between $x^{\prime}$ and $x_{1}$. Then, we can expand the expectation as follows:
$\mathbb{E}_{\mathbf{\Theta}^{\prime}}\Big[h\big(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}^{\prime}\big)\Big]=$ $\ \mathbb{E}_{\mathbf{\Theta}^{\prime}_{2:i}}\left[\mathbb{E}_{\Theta^{\prime}_{1}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}^{\prime}\right)\right]\right]$
$=$ $\ \Pr_{\Theta^{\prime}_{1}}\left[\Theta^{\prime}_{1}\in\Omega\right]~\mathbb{E}_{\mathbf{\Theta}^{\prime}_{2:i}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}^{\prime}\right)\mid\Theta^{\prime}_{1}\in\Omega\right]+$
$\ \Pr_{\Theta^{\prime}_{1}}\left[\Theta^{\prime}_{1}\notin\Omega\right]~\mathbb{E}_{\mathbf{\Theta}^{\prime}_{2:i}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}^{\prime}\right)\mid\Theta^{\prime}_{1}\notin\Omega\right]$
$=$ $\ \frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}~\left(1+\mathbb{E}_{\mathbf{\Theta}^{\prime}_{2:i}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}^{\prime}\right)\mid\Theta^{\prime}_{1}\in\Omega\right]\right)+$
$\ \frac{x_{i}-x_{1}}{x_{i}-x^{\prime}}~\left(1+\mathbb{E}_{\mathbf{\Theta}^{\prime}_{2:i}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}^{\prime}\right)\mid\Theta^{\prime}_{1}\notin\Omega\right]\right)\ .$
For $\Theta_{1}^{\prime}\in\Omega$, the tree growth process after the first split is the same as the case where $x^{\prime}$ is not added. Thus, we have
$\mathbb{E}_{\mathbf{\Theta}^{\prime}_{2:i}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}^{\prime}\right)\mid\Theta^{\prime}_{1}\in\Omega\right]=\mathbb{E}_{\mathbf{\Theta}_{1:i-1}}\left[h\left(x_{i};x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]\ .$
For $\Theta_{1}\notin\Omega$, the tree growth process is the same as the case where $x^{\prime}$ is added. Thus, we have
$\mathbb{E}_{\mathbf{\Theta}^{\prime}_{2:i}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}^{\prime}\right)\mid\Theta^{\prime}_{1}\notin\Omega\right]=\mathbb{E}_{\mathbf{\Theta}_{1:i-1}}\left[h\left(x_{i};x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]-1\ .$
@@
$\mathbb{E}_{\mathbf{\Theta}^{\prime}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}^{\prime}\right)\right]=$ $\ \frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}~\left(1+\mathbb{E}_{\mathbf{\Theta}^{\prime}_{2:i}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}^{\prime}\right)\mid\Theta^{\prime}_{1}\in\Omega\right]\right)+$
$\ \frac{x_{i}-x_{1}}{x_{i}-x^{\prime}}~\left(1+\mathbb{E}_{\mathbf{\Theta}^{\prime}_{2:i}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}^{\prime}\right)\mid\Theta^{\prime}_{1}\notin\Omega\right]\right)$
@@
$\mathbb{E}_{\mathbf{\Theta}^{\prime}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{n},\mathbf{\Theta}^{\prime}\right)\right]=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right]\ .$ (2)